### v1

# Explainable AI for Curie Temperature Prediction in Magnetic Materials

Abstract
We explore machine learning techniques for predicting Curie temperatures of magnetic materials using the NEMAD database. By augmenting the dataset with composition-based and domain-aware descriptors, we evaluate the performance of several machine learning models. We find that the Extra Trees Regressor delivers the best performance reaching an R2 score of up to 0.85 $\pm$ 0.01 (cross-validated) for a balanced dataset. We employ the k-means clustering algorithm to gain insights into the performance of chemically distinct material groups. Furthermore, we perform the SHAP analysis to identify key physicochemical drivers of Curie behavior, such as average atomic number and magnetic moment. By employing explainable AI techniques, this analysis offers insights into the model’s predictive behavior, thereby advancing scientific interpretability.
(a) Main dataset
(b) DS1 validation dataset
Figure 1: Predicted vs. true Curie temperatures using the ExtraTrees model for (a) the training/test dataset and (b) the DS1 validation dataset. A fairly strong alignment along the diagonal indicates the predictive performance of the model.
Figure 2: Feature importance graph for the Extra Trees Algorithm.
1  Introduction
Curie temperature (TC) is a critical property that defines the transition of a material from a ferromagnetic to a paramagnetic state. Predicting TC with high accuracy is essential for the discovery of magnetic materials for data storage, spintronics, and energy applications. Traditional approaches based on quantum mechanical computations or empirical models are often limited in scalability and accuracy. In recent years, machine learning (ML) has emerged as a promising alternative for property prediction across materials science domains [1, 2, 3, 4, 5].
Building on this momentum, several works have proposed the use of ML models trained on curated magnetic datasets. In particular, the recent study [6] introduced the NEMAD database, which aggregates experimentally measured magnetic transition temperatures and compositions. Similarly, the study by Belot et al. [7] utilized two of the largest available datasets of experimental Curie temperatures—comprising over 2,500 materials for training and more than 3,000 entries for validation—to compare machine learning strategies for predicting Curie temperature solely from chemical composition.
Our work is inspired by these prior efforts and aims to improve the predictive accuracy and gain insights into model interpretability. We develop a pipeline that starts from the NEMAD dataset, augments it with compositional and elemental features, and evaluates several ML models. A key contribution of our work is the integration of explainable AI (XAI) through SHAP (SHapley Additive exPlanations) analysis, which allows us to quantify how each input feature contributes to the model’s prediction. Moreover, we benchmark our models on external datasets from literature to demonstrate generalization.
2  Dataset Preparation and Model Evaluation
We begin our study with the NEMAD database [6], a publicly available dataset compiled using large language models, comprising 26,706 magnetic materials with records for chemical composition, phase transition temperatures (e.g., Curie or Néel temperatures), structural characteristics, and related magnetic properties. We pre-processed the database to remove ambiguous entries, duplicates, and entries with missing or non-numeric temperatures. We also standardized chemical formulas using the pymatgen library.
To generate descriptors, we first computed composition-based features such as average atomic number, atomic mass, group number, period, and electronegativity using pymatgen. We then added elemental property descriptors using the Magpie feature set via matminer, which includes statistics such as mean, range, and standard deviation of valence electron counts, ionic radii, and oxidation states. These features capture the diversity and bonding characteristics of constituent elements and have been shown to correlate with magnetic properties.
We evaluated several machine learning models for predicting Curie temperatures using the enriched dataset. Table 1 presents the comparative performance metrics in terms of Mean Absolute Error (MAE), Root Mean Squared Error (RMSE), and the coefficient of determination ($R^{2}$).
Table 1: Performance comparison of machine learning models for Curie temperature prediction after feature normalization where necessary.
Model MAE (K) RMSE (K) R2
Extra Trees 51 98 0.87
Random Forest 57 103 0.85
XGBoost 63 106 0.85
Neural Network 69 123 0.79
KNN 74 129 0.77
The Extra Trees algorithm achieved the best overall performance, with an $R^{2}$ value of 0.87, MAE of 54 K, and RMSE of 106 K. This performance surpasses that of the Random Forest algorithm and other models considered. With a balanced dataset we were able to achieve $R^{2}\sim 0.85$ with a standard deviation of $0.01$. The balanced dataset was obtained by randomly removing around 4% of samples with Curie temperatures less than 300 K [6].
The Extra Trees (Extremely Randomized Trees) algorithm employs an ensemble learning method similar to the Random Forests model, but with two notable distinctions:
  • It selects cut-points at random rather than optimizing splits.
  • It uses the entire training dataset without bootstrapping.
These differences introduce greater variance among the individual trees in the ensemble, which can improve generalization and reduce overfitting on noisy datasets. This may explain its superior performance in our analysis, where the feature space is high-dimensional and the temperature distribution is non-uniform.
The Random Forest model also performs well but slightly underperforms Extra Trees, likely due to its more deterministic node-splitting strategy, which may be less adaptive to irregularities in the dataset.
XGBoost, KNN, and neural networks were also tested. XGBoost shows competitive results but did not outperform tree-based ensembles. KNN and neural networks showed relatively weaker performance, potentially due to the complexity of the feature relationships and sensitivity to scaling.
We validate the performance of the Extra Trees model on the DS1 dataset [19]. Figure 1 presents the predicted versus actual Curie temperatures using the ExtraTrees model, both for the main dataset and the DS1 validation test dataset. In both subfigures, the data points cluster along the $y=x$ diagonal, indicating that the predicted and actual values agree fairly well ($R^{2}\sim 0.71$, MAE $\sim$ 91 K) which demonstrates the model’s generalizability.
(a) SHAP bar plot showing the average impact of each feature on model output.
(b) SHAP beeswarm plot showing feature influence per sample. Color indicates feature value.
Figure 3: SHAP analysis of the ExtraTrees model. (a) shows the average contribution of each feature across the dataset, while (b) reveals the variation of feature impacts for individual predictions.
3  SHAP Analysis
To interpret the predictions of our Extra Trees model in the context of Explainable AI, we use SHAP (SHapley Additive exPlanations). SHAP assigns each feature an importance value for a particular prediction, based on cooperative game theory. This allows us to quantify how much each input feature contributes to increasing or decreasing the predicted Curie temperature.
SHAP Bar Plot: Figure (3), left panel, shows the SHAP summary bar plot, which ranks the features by their average absolute SHAP values. This plot reveals that the Mean Magnetic Moment is the most influential feature. This descriptor captures the average magnetic moment of constituent atoms, which is strongly linked to Curie temperature, due to its role in long-range magnetic ordering. Features like Average Deviation of Number of Unfilled f-Orbitals and Mean Ground State Volume/Atom also show significant contributions, reflecting magnetic configuration and spatial arrangement of electrons.
Interestingly, elemental fractions like Fraction of Fe and Fraction of Mn have moderate importance, suggesting that while the presence of individual elements is informative, more complex electronic or structural descriptors may serve as stronger predictors of Curie temperature.
SHAP Beeswarm Plot: The SHAP beeswarm plot Figure (3), right panel, illustrates the distribution of SHAP values for each feature across all samples. Each point represents a single material, with color indicating the corresponding feature value. This visualization reveals not only which features are most influential, but also how they affect the model’s predictions. For example, higher values of the average magnetic moment of constituent atoms tend to increase the predicted Curie temperature, while lower values generally lead to decreased predictions.
We can see from the plot that the Mean Magnetic Moment, which represents the average magnetic moment of the ground-state structure, exhibits a strong positive influence when its value is high (shown in red), shifting SHAP values to the right. This suggests that materials with higher magnetic moments are strongly associated with higher Curie temperatures, which is physically intuitive.
In contrast, the average deviation of magnetic moment shows a somewhat different pattern. High values (in red) are generally associated with negative SHAP values, indicating a suppressive effect on Curie temperature. This indicates that when the distribution of atomic magnetic moments within a compound is more uneven, the model predicts a lower Curie temperature.
Additionally, the fraction of Mn feature, representing the proportion of manganese in the compound, shows a symmetrical SHAP value distribution with a subtle trend: higher Mn fractions (red) tend to push SHAP values upward. This behavior reinforces manganese’s well-established role in stabilizing magnetic ordering, which aligns with its frequent presence in high-$T_{C}$ magnetic materials.
These patterns not only validate domain expectations but also reveal hidden non-linearities that traditional feature importance metrics might overlook.
Figure 4: Elbow plot showing the inertia values for different numbers of clusters. The curve begins to flatten around $k=4$, suggesting a reasonable balance between variance capture and model complexity.
4  K-means Clustering Analysis
In this section, we present our implementation of the k-means clustering algorithm to identify subgroups within the dataset that may exhibit distinct magnetic behavior. To determine the optimal number of clusters, we employed the elbow method, as shown in Figure 4. The curve shows a sharp drop in inertia from $k=1$ to $k=3$, with diminishing returns beyond $k=4$. This led us to select $k=4$ for clustering.
We then applied the Extra Trees algorithm independently to each cluster and observed that Clusters 1 and 3 exhibited significantly lower predictive performance, with R2 values of approximately 0.58 and 0.38, respectively. Cluster 1 was found to be dominated by oxygen-rich compounds, particularly complex oxides, as illustrated in Figure 5. Examples of compounds in this cluster include, LaMnO3, SrFeO3, and CaMnO3. From a physical standpoint, such oxide systems differ substantially from the metallic systems that dominate other clusters. Their magnetism is governed by mechanisms like superexchange, strong electron correlation, and structural distortions (e.g., Jahn–Teller effects), which are not well captured by simple compositional descriptors. This explains the degradation in model performance within this cluster.
Cluster 3 yields the lowest $R^{2}$ compared to other clusters, with the highest MAE and RMSE values among all clusters. This underperformance likely arises from both its small size and its unique chemical composition. Unlike other clusters dominated by transition metal oxides or alloys, Cluster 3 contains a high proportion of carbon, and phosphorus-rich compounds, including intermetallics and carbides. These materials often exhibit complex bonding environments and non-collinear magnetic structures, which are challenging to model using composition-only descriptors. The limited number of training samples in this chemically diverse space further hinders generalization.
Therefore, the under performance of the model in Clusters 1 and 3 provides a strong motivation for incorporating structure-aware or quantum-level descriptors in future studies. Their distinct magnetic behaviors and modeling challenges highlight the limitations of current descriptor sets and the need for more physically grounded representations.
Consequently, excluding these clusters from our analysis improves overall model performance and this improves the value of $R^{2}$ to approximately 0.85. This decision is supported both statistically and physically and highlights the importance of domain knowledge when applying machine learning to datasets for magnetic materials. After removing Clusters 1 and 3 from the data, we obtain the following results from 5-fold cross validation:
Table 2: Cross-validation (5-fold) results for the best-performing model.
Metric Value Mean $\pm$ Std
MAE (K) 54 54 $\pm$ 2.5
RMSE (K) 105 105 $\pm$ 4.0
R2 0.85 0.85 $\pm$ 0.012
(a) Cluster-wise elemental composition percentages
(b) t-SNE projection with cluster labels
Figure 5: Figure 3(a) visualizes the dominant elements in each cluster, reinforcing chemically meaningful groupings. Figure 3(b) shows the t-SNE visualization for the dataset, where each point represents a material colored by its assigned cluster. The separation among clusters, particularly the distinct distribution of Cluster 1, suggests notable differences in structural or compositional properties that impact Curie temperature prediction.
t-SNE Visualization of Clusters
To further understand the separation and characteristics of clusters formed in the dataset, we employed t-SNE (t-distributed Stochastic Neighbor Embedding) to reduce the high-dimensional feature space to two dimensions. The resulting 2D visualization, shown in Figure 5, illustrates clear separation between chemically distinct groups.
Cluster 0 contains high concentrations of metallic elements such as iron (Fe), cobalt (Co), silicon (Si), and boron (B), reflecting typical magnetic intermetallics. Cluster 1 is dominated by complex oxides, rich in oxygen (O), manganese (Mn), lanthanum (La), and strontium (Sr)—a chemical signature common in perovskite and layered oxide systems. Cluster 2 is enriched in manganese (Mn), iron (Fe), gallium (Ga), and nickel (Ni), resembling Heusler-type alloys. Cluster 3 includes mainly carbon (C), iron (Fe), and Cobalt (Co), and represents simpler binary and ternary compounds.
These clusters not only form distinct regions in t-SNE space but also exhibit coherent chemical themes, validating the clustering strategy and emphasizing the influence of composition on Curie temperature behavior.
5  Conclusion
This study demonstrates that explainable machine learning can serve as a powerful tool for materials property prediction, particularly for estimating Curie temperatures using composition-based descriptors. We employed various machine learning models and found that, compared to other models, the optimized Extra Trees Regressor demonstrated robust predictive performance for Curie temperature estimation.
Furthermore, by leveraging explainable machine learning, we uncover feature-driven insights into the compositional factors that influence ferromagnetism in inorganic materials. The use of SHAP analysis allowed us to interpret model outputs in a physically meaningful way. Additionally, the application of clustering and t-SNE visualization revealed distinct patterns and subgroups within the dataset.
Our framework demonstrates that incorporating explainable AI not only improves transparency but also further strengthens the scientific value of machine learning models in materials science. This approach can be readily extended to other property prediction tasks across the domain.
6  Acknowledgments
We would like to thank Babar Shabbir for useful discussions.
The code and data for this project is available on github: https://github.com/dradeelajaib/Curie-Prediction